### 10 Mar 2024 v1

# Nonlinear Schrödinger equation in terms of elliptic and hyperelliptic $\sigma$ functions

Abstract.
It is known that the elliptic function solutions of the nonlinear Schrödinger equation are reduced to the algebraic differential relation in terms of the Weierstrass sigma function, ${\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha\frac{\partial}{\partial u}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\wp(v)+\lambda_{2}-\alpha^{2})\Psi}$, where $\Psi(u;v,t):={\mathrm{e}}^{\alpha u+{\mathfrak{i}}\beta t+c}$ ${\frac{{\mathrm{e}}^{-\zeta(v)u}\sigma(u+v)}{\sigma(u)\sigma(v)}}$, its dual $\Psi^{*}(u;v,t)$, and certain complex numbers $\alpha,\beta$. In this paper, we generalize the algebraic differential relation to those of genera two and three in terms of the hyperelliptic sigma functions.
Key words and phrases:
Nonlinear Schrödinger equation, hyperelliptic sigma functions
1991 Mathematics Subject Classification:
Primary 14H70, 14H05; Secondary 14H42, 14H45
1. Introduction
Our purpose of this work is to investigate the algebraic solutions of the Nonlinear Schrödinger (NLS) equation in terms of elliptic and hyperelliptic $\sigma$ functions.
As mentioned in Section 2, the elliptic solutions of the NLS equation are reduced to the algebraic differential relation [Ka],
(1.1) $\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha\frac{\partial}{\partial u}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\wp(v)+\lambda_{2}-\alpha^{2})\Psi.$
where
$\Psi(u;v,t):={\mathrm{e}}^{\alpha u+{\mathfrak{i}}\beta t+c}\frac{{\mathrm{e}}^{-\zeta(v)u}\sigma(u+v)}{\sigma(u)\sigma(v)},\quad\Psi^{*}(u):=-{\mathrm{e}}^{-\alpha u-{\mathfrak{i}}\beta t-c}\frac{{\mathrm{e}}^{-\zeta(v)u}\sigma(u-v)}{\sigma(u)\sigma(v)}$
for the elliptic curve $E:y^{2}=(x-e_{1})(x-e_{2})(x-e_{3})$, $c\in{\mathbb{C}}$, and the Weierstrass sigma functions of $E$. After we fix the point $v$ and $e_{i}$ so that $\Psi^{*}$ is the complex conjugate of $\Psi$ and $(\alpha,u)$ belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$, we find the elliptic solutions of the NLS equation. We call these conditions the reality conditions of the NLS equation. Determination of the conditions depends on the situations which we require, and is not simple because we tune the moduli parameter $e_{i}$ as in [Ka].
Construction of the algebraic solutions of the NLS equation was found by Previato 1985 [P] and is written in the book by Belokolos, Bobenko, Enolskii, Its and Matveev [BBEIM] precisely. On the other hand reevaluation of the theory of Kleinian sigma function, which was discovered by Klein have been performed by several authors [BEL1, BEL2, EEL, Ma6]. We have found the hyperelliptic solutions of the KdV equation [BEL1, Ma1], the MKdV equation [Ma1], the Toda lattice equation [KMP], the sine-Gordon equation [Ma3] in terms of the sigma functions and meromorphic functions of the symmetric product of algebraic curves. However the hyperelliptic solution of the NLS equation in terms of the Kleinian sigma functions has not been obtained yet. If we obtained it, it is easily to draw the numerical evaluations of the time development for a certain class of the curves [Ma5]. The approach is much easier than the approach by means of the the theta functions.
In this paper, besides the reality conditions, we show the generalization of the algebraic differential relation (1.1) of genera two and three in terms of the hyperelliptic sigma functions associated with the hyperelliptic solutions of the NLS equation. To obtain this, we express the Baker function in terms of the sigma function reported in [BEL1], though our Baker function slightly differs from that in [BEL1]. Using the Baker equation, we show the generalization. We note that the Baker equation, or the generalized Lamé equation, of the Baker function contains the potential $\wp_{gg}$, while the addition theorem of the sigma functions gives the meromorphic functions of the Jacobian $J_{X}$, which does not coincide with $\wp_{gg}$ itself except for $g=1$. Thus, the discrepancy prevents us from obtaining the generalization of (1.1). Even for genera two and three, we find the generalization of the algebraic differential relation 1.1) in this paper. In other words, the purpose of this paper shows how to overcome the discrepancy so that we obtain the generalization of (1.1) associated with the hyperelliptic solution of the NLS equation.
We note that in the NLS equation, we have crucial problems related to the reality conditions depending on the ”nonfocusing” and ”focusing” NLS equations. They correspond to the sign of the nonlinear terms in the NLS equation [P] and to the choice of $(\alpha,u)$ which belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$. However, we will not address these problems at this stage, as we plan to consider them in the next step.
The contents of this paper are as follows. Section 2 is devoted for the genus one case or the derivation of (1.1). Section 3 shows the preliminary of the hyperelliptic sigma function. Section 4 provides the addition formula related to this NLS equation. In Section 5, we show the Baker functions in terms of the hyperelliptic sigma functions. In Section 6, we generalize (1.1) for the genera two and three cases as by using the Baker function and the addition formulae in Theorems 6.2 and 6.4.
2. Genus one case
In this Section, we demonstrate how one gets an elliptic solution of the NLS equation (1.1) [Ka]. Let $E$ be an elliptic curve given by
$E:\ \frac{1}{4}{\widehat{y}}^{2}=y^{2}=x^{3}+\lambda_{2}x^{2}+\lambda_{1}x+\lambda_{0}$
[…]
where the $e$’s are complex numbers, $\lambda_{2}=-(e_{1}+e_{2}+e_{3})=0$ and ${\widehat{y}}=2y$.
The Weierstrass elliptic $\sigma$ function associated with the curve $E$ is connected with the Weierstrass Weierstrass-$\wp$ and $\zeta$ functions by
(2.2) $\wp(u)=-\frac{d^{2}}{du^{2}}\log\sigma(u),\quad\zeta(u)=\frac{d}{du}\log\sigma(u).$
Here the coordinate $u$ in the Jacobian $J_{E}$ associated with $E$ is given by
(2.3) $u=\int^{(x,y)}_{\infty}\frac{dx}{2y},$
with $x(u)=\wp(u),~{}{\widehat{y}}(u)=\wp^{\prime}(u)$ and $\infty$ is the infinity point of $E$. The Jacobian is given by $J_{E}={\mathbb{C}}/(2{\mathbb{Z}}\omega^{\prime}+2{\mathbb{Z}}\omega^{\prime\prime})$ using the double period $(2\omega^{\prime},2\omega^{\prime\prime})$.
The key to obtain a Weierstrass-$\wp$ function solution of the NLS equation is the addition formula,
(2.4) $\wp(u)-\wp(v)=-\frac{\sigma(u+v)\sigma(u-v)}{[\sigma(v)\sigma(u)]^{2}}.$
Let $(\wp(u),\wp^{\prime}(u))=(x_{1},2y_{1})$ and $(\wp(v),\wp^{\prime}(v))=(x_{2},2y_{2})$. By using ${\frac{d}{du}=2y_{1}\frac{d}{dx_{1}}}$ and ${\frac{d}{dv}=2y_{2}\frac{d}{dx_{2}}}$, we find that (2.4) leads the following relations:
(2.5) $\zeta(u+v)+\zeta(u-v)-2\zeta(u)=-\frac{2y_{1}}{x_{2}-x_{1}},$
(2.6) $\zeta(u+v)-\zeta(u)-\zeta(v)=\frac{y_{2}-y_{1}}{x_{2}-x_{1}},$
(2.7) $\wp(u+v)+\wp(u-v)=\frac{2f_{\Omega}(x_{1},x_{2})}{(x_{2}-x_{1})^{2}},$
(2.8) $\wp(u-v)=\frac{f_{\Omega}(x_{1},x_{2})+2y_{1}y_{2}}{(x_{2}-x_{1})^{2}},$
(2.9) $\wp(u+v)+\wp(u)+\wp(v)=\left[\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\right]^{2}-\lambda_{2}.$
where
(2.10) $f_{\Omega}(x_{1},x_{2})=x_{1}x_{2}(x_{1}+x_{2}+2\lambda_{2})+\lambda_{1}(x_{1}+x_{2})+\lambda_{0}.$
By using (2.5)-(2.9) and by defining $\psi(u;v)=\displaystyle{\frac{{\mathrm{e}}^{-\zeta(v)u}\sigma(u+v)}{\sigma(u)\sigma(v)}}$, we have the relation,
(2.11) $\left[\frac{d^{2}}{du^{2}}-2\wp(u)\right]\psi(u;v)=(\wp(v)+\lambda_{2})\psi(u;v),$
which is known as the Lamé equation. $\psi$ is known as the Baker function, or the Baker-Akhiezer function [BEL1, Ma6].
We also have ${\displaystyle\frac{d\psi(u;v)}{du}}=(\zeta(u+v)-\zeta(u)-\zeta(v))\psi(u;v)$, or
(2.12) $\left[\frac{d}{du}-(\zeta(u+v)-\zeta(u))\right]\psi(u;v)=\zeta(v)\psi(u;v).$
Further, by the Miura relation, $2x+\lambda_{2}+{\widetilde{x}}=\mu^{2}+\mu_{u}$ for ${\mu:=\frac{d}{du}\log\psi(u;v)=\frac{y-{\widetilde{y}}}{x-{\widetilde{x}}}}$, we have the static MKdV equation,
(2.13) $02(\lambda_{2}+3{\widetilde{x}})\mu_{u}+6\mu^{2}\mu_{u}+\mu_{uuu}=$ 0 .
We introduce
$\Psi(u;v,t):={\mathrm{e}}^{\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u):=-{\mathrm{e}}^{-\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v),$
for $\alpha,\beta,c\in{\mathbb{C}}$. Then the addition theorem (2.4) leads the relation,
(2.14) $\left[{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha\frac{\partial}{\partial u}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta-\wp(v)+\lambda_{2}-\alpha^{2})\Psi,$
which is the same as (1.1).
3. Hyperelliptic curve $X$ and sigma functions
In this Section, we give a background information on the hyperelliptic $\theta$ functions and the $\sigma$ functions as a generalization of the Weierstrass elliptic $\sigma$ function.
3.1. Geometrical setting of hyperelliptic curves
Let $X$ be a hyperelliptic curve defined by
$X~{}:~{}y^{2}=f(x):=x^{2g+1}+\lambda_{2g}x^{2g}+\cdots+\lambda_{0}$
together with a smooth point $\infty$ at infinity. Here $\lambda$’s are complex numbers. The affine ring related to $X$ is denoted by $R_{X}:={\mathbb{C}}[x,y]/(y^{2}-f(x))$. We fix a basis of holomorphic one-forms
${\nu^{\mathrm{I}}_{i}}=\frac{x^{i-1}dx}{2y}\qquad(i=1,\ldots,g),$
and the homology basis for the curve $X$,
$\mathrm{H}_{1}(X,\mathbb{Z})=\bigoplus_{j=1}^{g}\mathbb{Z}\alpha_{j}\oplus\bigoplus_{j=1}^{g}\mathbb{Z}\beta_{j},$
so that their intersections are given by $[\alpha_{i},\alpha_{j}]=0$, $[\beta_{i},\beta_{j}]=0$ and $[\alpha_{i},\beta_{j}]=-[\beta_{i},\alpha_{j}]\delta_{ij}$. We take the half-period matrix $\omega=\left[\begin{matrix}\omega^{\prime}\\ \omega^{\prime\prime}\end{matrix}\right]$ of $X$ with respect to the given basis where
$\omega^{\prime}=\frac{1}{2}\left[\oint_{\alpha_{j}}{\nu^{\mathrm{I}}_{i}}\right],\quad\omega^{\prime\prime}=\frac{1}{2}\left[\oint_{\beta_{j}}{\nu^{\mathrm{I}}_{i}}\right],$
The period matrices $2\omega^{\prime}$ and $2\omega^{\prime\prime}$ form the lattice $\Gamma_{X}$ in ${\mathbb{C}}^{g}$ as a ${\mathbb{Z}}$-module. The Jacobian variety of $X$ is denoted by $J_{X}$, $\kappa_{J}:{\mathbb{C}}^{g}\to J_{X}={\mathbb{C}}^{g}/\Gamma_{X}$. The abelian covering of $X$ generated by the path space of $X$ with the basepoint $\infty\in X$ is denoted by ${\widetilde{X}}$, $\kappa_{X}:{\widetilde{X}}\to X$ $\kappa_{X}(\gamma_{P,\infty})=P$, where $\gamma_{P,\infty}$ is a path from $\infty$ to $P\in X$. For a non-negative integer $k$, we define the Abelian integral ${\widetilde{w}}$ from $k$-th symmetric product $S^{k}{\widetilde{X}}$ of ${\widetilde{X}}$ to ${\mathbb{C}}^{g}$ and the Abel-Jacobi map $w$ from $S^{k}X$ by
${\widetilde{w}}:S^{k}{\widetilde{X}}\to{\mathbb{C}},\quad{\widetilde{w}}(\gamma_{1},\ldots,\gamma_{k})=\sum_{i=1}^{k}\int_{\gamma_{i}}\begin{pmatrix}{\nu^{\mathrm{I}}_{1}}\\ \vdots\\ {\nu^{\mathrm{I}}_{g}}\end{pmatrix},$
$w:S^{k}X\to J_{X},\quad w((x_{1},y_{1}),\cdots,(x_{k},y_{k}))=\sum_{i=1}^{k}\int_{\infty}^{(x_{i},y_{i})}\begin{pmatrix}{\nu^{\mathrm{I}}_{1}}\\ \vdots\\ {\nu^{\mathrm{I}}_{g}}\end{pmatrix}\hbox{\rm mod}\ {\Lambda}.$
The image of $w$ is denoted by $W_{X}^{k}=w(S^{k}X)$, i.e., $J_{X}=W_{X}^{g}$. Further we introduce an injection $\iota_{X}:X\to{\widetilde{X}}$ and will fix it. We find that $w=\kappa_{J}\circ{\widetilde{w}}\circ\iota_{X}$.
Definition 3.1.
Let $(x_{i},y_{i})_{i=1,\ldots,g}\in S^{g}X$.
  1. (1)
    We define the polynomials associated with $F(x)=(x-x_{1})\cdots(x-x_{g})$ by
    (3.1) $\pi_{i}(x):=\frac{F(x)}{x-x_{i}}=\chi_{i,g-1}x^{g-1}+\chi_{i,g-2}x^{g-2}+\cdots+\chi_{i,1}+\chi_{i,0},$
  2. (2)
    We define $g\times g$ matrices as follows.
    ${\mathcal{X}}_{g}:={\small{\begin{pmatrix}\chi_{1,0}&\chi_{1,1}&\cdots&\chi_{1,g-1}\\ \chi_{2,0}&\chi_{2,1}&\cdots&\chi_{2,g-1}\\ \vdots&\vdots&\ddots&\vdots\\ \chi_{g,0}&\chi_{g,1}&\cdots&\chi_{g,g-1}\end{pmatrix}}},\quad{\mathcal{Y}}_{g}:={\small{\begin{pmatrix}y_{1}&\ &\ &\ \\ \ &y_{2}&\ &\ \\ \ &\ &\ddots&\ \\ \ &\ &\ &y_{g}\end{pmatrix}}}=\frac{1}{2}{\mathcal{H}}_{g},$
    ${\mathcal{F}}_{g}^{\prime}:={\small{\begin{pmatrix}F^{\prime}(x_{1})&&&\\ &F^{\prime}(x_{2})&&\\ &&\ddots&\\ &&&F^{\prime}(x_{g})\end{pmatrix}}},{\mathcal{V}}_{g}:={\small{\begin{pmatrix}1&1&\cdots&1\\ x_{1}&x_{2}&\cdots&x_{g}\\ x_{1}^{2}&x_{2}^{2}&\cdots&x_{g}^{2}\\ \cdot&\cdot&&\cdot\\ x_{1}^{g-1}&x_{2}^{g-1}&\cdots&x_{g}^{g-1}\end{pmatrix}}},$
    ${\mathcal{E}}_{g}:={\small{\begin{pmatrix}1&\ &\ &\ &\ \\ \varepsilon_{g-1}&1&\ &\text{\Large 0}&\ \\ \varepsilon_{g-2}&\varepsilon_{g-1}&1&&\ \\ \vdots&\vdots&\vdots&\ddots&\\ \varepsilon_{1}&\varepsilon_{2}&\cdots&\varepsilon_{g-1}&1\\ \end{pmatrix}}},\quad{\mathcal{K}}_{g}:={\small{\begin{pmatrix}x_{1}^{g-1}&x_{1}^{g-2}&\cdots&1\\ x_{2}^{g-1}&x_{2}^{g-2}&\cdots&1\\ \vdots&\vdots&\ddots&\vdots\\ x_{g}^{g-1}&x_{g}^{g-2}&\cdots&1\end{pmatrix}}}.$
    where $F(x)=x^{g}+\varepsilon_{g-1}x^{g-1}+\cdots\varepsilon_{1}x+\varepsilon_{0}$ and $F^{\prime}(x):=dF(x)/dx$.
Lemma 3.2.
Let $u\in{\widetilde{w}}(\iota_{X}((x_{1},y_{1}),\ldots,(x_{g},y_{g})))$.
  1. (1)
    The inverse matrix of ${\mathcal{X}}_{g}$ is given by ${\mathcal{X}}_{g}^{-1}={\mathcal{F}}_{g}^{\prime-1}{\mathcal{V}}_{g}$.
  2. (2)
    For $\partial_{u_{i}}:=\partial/\partial{u_{i}}$, $\partial_{x_{i}}:=\partial/\partial{x_{i}}$, and $\partial_{u_{i}}^{(r)}:=\partial/\partial{u_{i}^{(r)}}$, we have
    ${\small{\begin{pmatrix}\partial_{u_{1}}\\ \partial_{u_{2}}\\ \vdots\\ \partial_{u_{g}}\end{pmatrix}}}=2{\mathcal{Y}}_{g}{\mathcal{F}}_{g}^{\prime-1}{\mathcal{X}}_{g}{\small{\begin{pmatrix}\partial_{x_{1}}\\ \partial_{x_{2}}\\ \vdots\\ \partial_{x_{g}}\end{pmatrix},}}\quad\frac{\partial x_{i}}{\partial u_{r}}=\frac{2y_{i}}{F^{\prime}(x_{i})}\chi_{i,r-1},\quad\frac{\partial y_{i}}{\partial u_{r}}=\frac{f^{\prime}(x_{i})}{F^{\prime}(x_{i})}\chi_{i,r-1},$
    (3.2) $\frac{\partial}{\partial u_{g}}=\sum_{i=1}^{g}\frac{2y_{i}}{F^{\prime}(x_{i})}\frac{\partial}{\partial x_{i}},\quad\frac{\partial}{\partial u_{g-1}}=\sum_{i=1}^{g}\frac{2y_{i}\chi_{i,g-1}}{F^{\prime}(x_{i})}\frac{\partial}{\partial x_{i}}.$
3.2. Sigma function and its derivatives
We define differentials of the second kind,
${\nu^{\mathrm{II}}_{j}}={\displaystyle\frac{1}{2y}}\sum_{k=j}^{2g-j}(k+1-j)\lambda_{k+1+j}x^{k}dx,\quad(j=1,\cdots,g)$
and complete hyperelliptic integrals of the second kind
$\eta^{\prime}=\frac{1}{2}\left[\oint_{\alpha_{j}}{\nu^{\mathrm{II}}_{i}}\right],\quad\eta^{\prime\prime}=\frac{1}{2}\left[\oint_{\beta_{j}}{\nu^{\mathrm{II}}_{i}}\right].$
For this basis of the $2g$-dimensional space of meromorphic differentials, the half-periods $\omega^{\prime},\omega^{\prime\prime},\eta^{\prime},\eta^{\prime\prime}$ satisfy the generalized Legendre relation
(3.3) ${\mathfrak{M}}\left(\begin{array}[]{cc}0&-1_{g}\\ 1_{g}&0\end{array}\right){\mathfrak{M}}^{T}=\frac{{\mathfrak{i}}\pi}{2}\left(\begin{array}[]{cc}0&-1_{g}\\ 1_{g}&0\end{array}\right),$
where $\mathfrak{M}=\left(\begin{array}[]{cc}\omega^{\prime}&\omega^{\prime\prime}\\ \eta^{\prime}&\eta^{\prime\prime}\end{array}\right)$. Let $\tau={\omega^{\prime}}^{-1}\omega^{\prime\prime}$. The theta function on $\mathbb{C}^{g}$ with modulus $\tau$ and characteristics $\tau a+b$ for $a,b\in{\mathbb{C}}^{g}$ is given by
$\theta\negthinspace\left[\begin{matrix}a\\ b\end{matrix}\right](z;\tau)=\sum_{n\in\mathbb{Z}^{g}}\exp\left[2\pi i\left\{{\displaystyle\frac{1}{2}}\ ^{t}\,\negthinspace(n+a)\tau(n+a)+\ ^{t}\,\negthinspace(n+a)(z+b)\right\}\right].$
The $\sigma$-function ([B1], p.336,[BEL1]), an analytic function on the space ${\mathbb{C}}^{g}$ and a theta series having modular invariance of a given weight with respect to $\mathfrak{M}$, is given by the formula
$\sigma(u)=\gamma_{0}\,\mathrm{exp}\left\{-{\displaystyle\frac{1}{2}}\ ^{t}u\eta^{\prime}{{\omega}^{\prime}}^{-1}u\right\}\theta\negthinspace\left[\begin{matrix}\delta^{\prime\prime}\\ \delta^{\prime}\end{matrix}\right]\left(\frac{1}{2}{{\omega}^{\prime}}^{-1}u\,;\,\mathbb{T}\right),$
where $\delta^{\prime}$ and $\delta^{\prime\prime}$ are half-integer characteristics giving the vector of Riemann constants with basepoint at $\infty$ and $\gamma_{0}$ is a certain non-zero constant. The $\sigma$-function vanishes only on $\kappa_{J}^{-1}(W_{X}^{g-1})$ (see for example [B1], p. 252). The hyperelliptic Weierstrass-$\wp$ and $\zeta$ functions are defined by
$\wp_{ij}=-\frac{\partial^{2}}{\partial u_{i}\partial u_{j}}\log\sigma(u),\quad\zeta_{i}=\frac{\partial}{\partial u_{i}}\log\sigma(u).$
As the Jacobi inversion formula,
$F(x)=x^{g}-\sum_{i=1}^{g}\wp_{gi}(u)x^{i-1}.$
Let $\{\phi_{i}(x,y)\}$ be an ordered set of ${}{\mathbb{C}}\cup\{\infty\}$-valued functions over $X$ defined by
(3.7) $\phi_{i}(x,y)=\left\{\begin{array}[]{llll}x^{i}&\mbox{for }i\leq g,\\ x^{\lfloor(i-g)/2\rfloor+g}&\mbox{for }i>g~{}{\rm and}~{}i-g\mbox{ even,}\\ x^{\lfloor(i-g)/2\rfloor}y&\mbox{for }i>g~{}{\rm and}~{}i-g\mbox{ odd.}\\ \end{array}\right.$
Here we note that $\{\phi_{i}(x,y)\}$ is a set of the bases of $R_{X}$ as the ${\mathbb{C}}$ vector space.
Following [O], we introduce a multi-index $\natural^{n}$. For $n$ with $1\leq n<g$, we let $\natural^{n}$ be the set of positive integers $i$ such that $n+1\leq i\leq g$ with $i\equiv n+1$ mod $2$. Namely,
$\natural^{n}=\left\{\begin{array}[]{lll}n+1,n+3,\cdots,g-1&\text{for}\;g-n\equiv 0\;\mathrm{mod}\;2\\ n,n+2,\cdots,g&\text{for}\;g-n\equiv 1\;\mathrm{mod}\;2\\ \end{array}\right.$
and partial derivative over the multi-index $\natural^{n}$
$\sigma_{\natural^{n}}=\bigg{(}\prod_{i\in\natural^{n}}\frac{\partial}{\partial u_{i}}\bigg{)}\sigma(u)$
For $n\geq g$, we define $\natural^{n}$ as empty and $\sigma_{\natural^{n}}$ as $\sigma$ itself.
Table 1
genus $\sigma_{\natural^{1}}$ $\sigma_{\natural^{2}}$ $\sigma_{\natural^{3}}$ $\sigma_{\natural^{4}}$ $\sigma_{\natural^{5}}$ $\sigma_{\natural^{6}}$ $\sigma_{\natural^{7}}$ $\sigma_{\natural^{8}}$cdots $1$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$cdots $2$ $\sigma_{2}$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$cdots $3$ $\sigma_{2}$ $\sigma_{3}$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$cdots $4$ $\sigma_{24}$ $\sigma_{3}$ $\sigma_{4}$ $\sigma$ $\sigma$ $\sigma$ $\sigma$ $\sigma$cdots $5$ $\sigma_{24}$ $\sigma_{35}$ $\sigma_{4}$ $\sigma_{5}$ $\sigma$ $\sigma$ $\sigma$ $\sigma$cdots $6$ $\sigma_{246}$ $\sigma_{35}$ $\sigma_{46}$ $\sigma_{5}$ $\sigma_{6}$ $\sigma$ $\sigma$ $\sigma$cdots $7$ $\sigma_{246}$ $\sigma_{357}$ $\sigma_{46}$ $\sigma_{57}$ $\sigma_{6}$ $\sigma_{7}$ $\sigma$ $\sigma$cdots $8$ $\sigma_{2468}$ $\sigma_{357}$ $\sigma_{468}$ $\sigma_{57}$ $\sigma_{68}$ $\sigma_{7}$ $\sigma_{8}$ $\sigma$cdots $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\ddots$
For $u\in{\mathbb{C}}^{g}$, we denote by $u^{\prime}$ and $u^{\prime\prime}$ the unique vectors in $\mathbb{R}^{g}$ such that
$u=2\,{}^{t}\omega^{\prime}u^{\prime}+2\,{}^{t}\omega^{\prime\prime}u^{\prime\prime}.$
We define
$L(u,v)$ $={}^{t}{u}(2\,{}^{t}\eta^{\prime}v^{\prime}+2\,{}^{t}\eta^{\prime\prime}v^{\prime\prime}),$
$\chi(\ell)$ $=\exp\big{\{}2\pi i\big{(}{}^{t}{\ell^{\prime}}\delta^{\prime\prime}-{}^{t}{\ell^{\prime\prime}}\delta^{\prime}+\tfrac{1}{2}{}^{t}{\ell^{\prime}}\ell^{\prime\prime}\big{)}\big{\}}\ (\in\{1,-1\})$
for $u$, $v\in{\mathbb{C}}^{g}$ and for $\ell$ ($=2\,{}^{t}\omega^{\prime}\ell^{\prime}+2\,{}^{t}\omega^{\prime\prime}\ell^{\prime\prime}$) $\in\Lambda$. Then $\sigma_{\natural^{n}}(u)$ for $u\in\kappa_{J}^{-1}(W_{X}^{n})$ satisfies the translational relation ([O], Lemma 7.3):
(3.8) $\sigma_{\natural^{n}}(u+\ell)=\chi(\ell)\sigma_{\natural^{n}}(u)\exp L(u+\tfrac{1}{2}\ell,\ell)\ \ \hbox{for $u\in\kappa_{J}^{-1}(W_{X}^{n})$}.$
Then we also have the translational formulas of Weierstrass-$\wp$ and $\zeta$ for $u\in{\mathbb{C}}^{g}$,
(3.9) $\wp_{ij}(u+\ell)=\wp_{ij}(u),\quad\zeta_{i}(u+\ell)=\zeta_{i}(u)+\sum_{j=1}^{g}(\eta_{ij}^{\prime}\ell_{j}^{\prime}+\eta_{ij}^{\prime\prime}\ell_{j}^{\prime\prime}).$
Further for $n\leqq g$, we note that $\sigma_{\natural^{n}}(-u)=(-1)^{ng+\frac{1}{2}n(n-1)}\sigma_{\natural^{n}}(u)$ for $u\in\kappa_{J}^{-1}(W_{X}^{n})$, especially,
(3.10) $\left\{\begin{array}[]{llll}\sigma_{\flat}(-u)=-\sigma_{\flat}(u)&\hbox{for $u\in\kappa_{J}^{-1}(W_{X}^{2})$}\\[6.45831pt] \sigma_{\sharp}(-u)=(-1)^{g}\sigma_{\sharp}(u)&\hbox{for $u\in\kappa_{J}^{-1}(W_{X}^{1})$}\end{array}\right.$
by Proposition 7.5 in [O].
4. The addition formulae
In this Section, we give the addition formulae of the hyperelliptic $\sigma$ functions which are the generalization of the addition formula (2.4). Those then give the key to construct the hyperelliptic solution of the Toda lattice.
4.1. The algebraic addition formula
Let us first recall the generalization of the addition formula (2.4):
Definition 4.1.
For a positive integer $n\geq 1$ and $(x_{1},y_{1}),\cdots,(x_{n},y_{n})$ in $X$, we define the Frobenius-Stickelberger determinant [KMP, Ma6],
$\begin{split}&\Psi_{n}((x_{1},y_{1}),\cdots,(x_{n},y_{n}))\\ &:=\left|\begin{matrix}1&\phi_{1}(x_{1},y_{1})&\cdots&\phi_{n-2}(x_{1},y_{1})&\phi_{n-1}(x_{1},y_{1})\\ 1&\phi_{1}(x_{2},y_{2})&\cdots&\phi_{n-2}(x_{2},y_{2})&\phi_{n-1}(x_{2},y_{2})\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 1&\phi_{1}(x_{n-1},y_{n-1})&\cdots&\phi_{n-2}(x_{n-1},y_{n-1})&\phi_{n-1}(x_{n-1},y_{n-1})\\ 1&\phi_{1}(x_{n},y_{n})&\cdots&\phi_{n-2}(x_{n},y_{n})&\phi_{n-1}(x_{n},y_{n})\\ \end{matrix}\right|,\end{split}$
where $\phi_{i}(x_{j},y_{j})$’s are the monomials defined in (3.7).
Let us first note that the curve $X$ has the algebraic addition structure [KMP, Ma6]:
Definition 4.2.
For given $P_{1},\cdots,P_{n}\in X$, we define
$\mu_{n}(P;P_{1},\cdots,P_{n})=\lim_{Q_{i}\to P_{i}}\frac{\Psi_{n+1}(P,Q_{1},\cdots,Q_{n})}{\Psi_{n}(Q_{1},\cdots,Q_{n})}$
for distinct $Q_{i}$’s.
Proposition 4.3.
For given $P_{1},\cdots,P_{n}\in X$, we have $Q_{i},\cdots,Q_{\ell}$ with $\ell=g$ for $n\geq g$ and $\ell=n$ otherwise, such that
$P_{1}+P_{2}+\cdots+P_{n}+Q_{1}+Q_{2}+\cdots+Q_{\ell}-(n+\ell)\infty\sim 0$
as a non trivial zero of $\mu_{n}(P;P_{1},\cdots,P_{n})$. For each $Q_{i}=(x_{i},y_{i})$, by letting $-Q_{i}=(x_{i},-y_{i})$, we have the addition property,
$P_{1}+P_{2}+\cdots+P_{n}-n\infty\sim(-Q_{1})+(-Q_{2})+\cdots+(-Q_{\ell})-(-\ell)\infty.$
Remark 4.4.
We should note that the Hyperelliptic involution $\iota:(x,y)\mapsto(x,-y)$, which induces $[-1]$-action on $J_{X}$ such that $u\mapsto-u$.
In this sense, since we fix the base point $\infty$, we use the notations, $n(x,y)$ for a point $(x,y)\equiv P\in X$, the addition $(x,y)+(x^{\prime},y^{\prime})$, and equality “$=$”; for example $(x,y)+(x^{\prime},y^{\prime})\neq(x+x^{\prime},y+y^{\prime})$ and $2(x,y)\neq(2x,2y)$ in general. Further we note that for every $P\in X$, $2P=-\sum_{i=1}^{2g-1}Q_{i}$ for points $Q_{i}$’s.
4.2. The analytic addition formula
We have the addition formula for the hyperelliptic $\sigma$ functions as the generalization of (2.4). (Theorem 5.1 in [EEMOP]):
Theorem 4.5.
Assume that $(m,n)$ is a pair of positive integers. Let $(x_{i},y_{i})$ $(i=1,\cdots,m)$, $(x^{\prime}_{j},y^{\prime}_{j})$ $(j=1,\cdots,n)$ in $X$ and $u\in\kappa_{J}^{-1}(W_{X}^{m})$, $v\in\kappa_{J}^{-1}(W_{X}^{n})$ be points such that $u=w((x_{1},y_{1}),\cdots,(x_{m},y_{m}))$ and $v=w((x_{1}^{\prime},y_{1}^{\prime}),\cdots,(x_{n}^{\prime},y_{n}^{\prime}))$. Then the following relation holds :
(4.1) $\begin{split}&\frac{\sigma_{\natural^{m+n}}(u+v)\sigma_{\natural^{m+n}}(u-v)}{\sigma_{\natural^{m}}(u)^{2}\sigma_{\natural^{n}}(v)^{2}}\\ &=\delta(g,m,n)\frac{\prod_{i=0}^{1}\Psi_{m+n}((x_{1},y_{1}),\cdots,(x_{m},y_{m}),(x_{1}^{\prime},(-1)^{i}y_{1}^{\prime}),\cdots,(x_{n}^{\prime},(-1)^{i}y_{n}^{\prime}))}{(\Psi_{m}((x_{1},y_{1})\cdots,(x_{m},y_{m}))\Psi_{n}((x_{1}^{\prime},y_{1}^{\prime})\cdots,(x_{n}^{\prime},y_{n}^{\prime})))^{2}}\\ &\times\prod_{i=1}^{m}\prod_{j=1}^{n}\frac{1}{\Psi_{2}((x_{i},y_{i}),(x_{j}^{\prime},y^{\prime}_{j}))},\end{split}$
where $\delta(g,m,n)=(-1)^{gn+\frac{1}{2}n(n-1)+mn}$.
Theorem 4.5 with $m=g$ and $n=1$ leads to the following Corollary:
Belokolos, Enolskii, and Salerno found the following relation [BES, Theorem 3.2];
Corollary 4.6.
Let $(x_{i},y_{i})\in X$ $(i=1,\cdots,g)$, $({\widetilde{x}},{\widetilde{y}}):=(x^{\prime}_{1},y^{\prime}_{1})\in X$, $u\in{\mathbb{C}}^{g}$ and $v\in\kappa_{J}^{-1}(W_{X}^{1})$ such that $u=w((x_{1},y_{1}),\cdots,(x_{g},y_{g}))$ modulo $\Gamma_{X}$ and $v=w(({\widetilde{x}},{\widetilde{y}}))$ modulo $\Gamma_{X}$. Then we have
(4.2) $\begin{split}\frac{\sigma(u+v^{[1]})\sigma(u-v^{[1]})}{\sigma(u)^{2}\sigma_{\flat}(v^{[1]})^{2}}&={{\widetilde{x}}}^{g}-\sum_{i=1}^{g}\wp_{gj}(u){x^{\prime}_{1}}^{i+j-2}\\ &=F({\widetilde{x}})=({\widetilde{x}}-x_{1})({\widetilde{x}}-x_{2})\cdots({\widetilde{x}}-x_{g}).\end{split}$
Proof.
Direct computation shows it.
Noting the difference in the differential operators in Jacobian $J_{X}$, we generalize (2.5)-(2.9).
Lemma 4.7.
For $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\ldots,\iota_{X}(x_{g},y_{g}))$ and $v={\widetilde{w}}(\iota_{X}(x,y))$ in the case of $k=g$ of Corollary 4.6, the following hold.
  1. (1)
    ${\sum_{i=1}^{g}e^{i-1}\left(\zeta_{i}(u+v)+\zeta_{i}(u-v)-2\zeta_{i}(u)\right)=-\sum_{\ell=1}^{g}\frac{2y_{\ell}F(e)}{(e-x_{\ell})(x-x_{\ell})F^{\prime}(x_{\ell})}=-F(e)\Delta_{x,e},}$
  2. (2)
    ${\zeta_{i}(u+v)+\zeta_{i}(u-v)-2\zeta_{i}(u)=-\sum_{\ell=1}^{g}\frac{2y_{\ell}\chi_{\ell,i-1}}{(x-x_{\ell})F^{\prime}(x_{\ell})},}$
  3. (3)
    ${\sum_{i=1}^{g}x_{r}^{i-1}\left(\zeta_{i}(u+v)+\zeta_{i}(u-v)-2\zeta_{i}(u)\right)=-\frac{2y_{r}}{(x-x_{r})},}$
  4. (4)
    ${\sum_{i=1}^{g}x^{i-1}\left(\zeta_{i}(u+v)-\zeta_{i}(u-v)-2\zeta_{i}^{\natural_{1}}(v)\right)=\frac{2yF^{\prime}(x)}{F(x)}=\sum_{\ell=1}^{g}\frac{2y}{(x-x_{\ell})},}$
  5. (5)
    ${\sum_{i,j}^{g}x^{i-1}x_{r}^{i-1}\left(\wp_{ij}(u+v)-\wp_{ij}(u-v)\right)=-\frac{4y_{r}y}{(x-x_{r})^{2}}.}$
where ${\zeta_{i}^{\natural_{j}}(v):=\left[\frac{\partial}{\partial u_{i}}\log\sigma_{\natural_{j}}(u)\right]_{u=v}}$, and $e$ is a generic parameter.
Proof.
Lemma 3.2 and (3.1) show 1. Noting $\partial/\partial x=\sum_{i}(2y/x^{i-1})\partial/\partial v_{i}$, we have 3, 4, and 5.
Proposition 4.8.
Let $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\ldots,\iota_{X}(x_{g},y_{g}))$, and $v={\widetilde{w}}(\iota_{X}(x,y))$a.
(4.3) $\wp_{gg}(u+v)+\wp_{gg}(u-v)+2\wp_{gg}(u)=2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+2\frac{f(x)}{F(x)^{2}}-2\lambda_{2g}-2x.$
Proof.
We consider the differential of 4.7 2 with respect to $u$. We interpret the differential (3.2) with respect to $u$ into differentials with respect to $x_{r}$, and then we have
$\begin{split}\wp_{gg}(u+v)&+\wp_{gg}(u-v)-2\wp_{gg}(u)=\sum_{\ell}\frac{2y_{\ell}}{F^{\prime}(x_{\ell})}\frac{\partial}{\partial x_{\ell}}\sum_{r}\frac{2y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\\ &=2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+\sum_{r}\frac{2}{F^{\prime}(x_{r})}\frac{\partial}{\partial x}\left[\frac{f(x)}{(x-x)F^{\prime}(x)}\right]\Bigr{|}_{x=x_{r}}\\ &=2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+2\left[\frac{f(x)}{F(x)^{2}}-\lambda_{2g}-x-2\wp_{gg}(u)\right].\\ \end{split}$
The second equality is asserted by Lemma 4.9. Lemma 4.10 shows the final step.
Lemma 4.9.
${\frac{\partial}{\partial x_{k}}F^{\prime}(x_{k})=\frac{1}{2}\left[\frac{\partial^{2}}{\partial x^{2}}F(x)\right]_{x=x_{k}}}$.
Lemma 4.10.
For $({\widetilde{x}},{\widetilde{y}})\in X$ and $F(x)$, we have
${\sum_{k=1}^{g}\frac{1}{F^{\prime}(x_{k})}\left[\frac{\partial}{\partial x}\left(\frac{f(x)}{(x-{\widetilde{x}})F^{\prime}(x)}\right)\right]_{x=x_{k}}=-\frac{f({\widetilde{x}})}{F({\widetilde{x}})^{2}}+\lambda_{2g}+b_{r}+2\wp_{gg}.}$
Proof.
The relation is proved by the integral.
  1. (1)
    Let us consider the polygon expansion of the curve $X$ by $D_{X}$ associated with the standard homology basis. Let its boundary be denoted by $\partial D_{X}$.
[…]
  2. (2)
    The divisor of integrand in (4.4) is
    (4.5) $\left(\frac{f(x)}{(x-b_{r})F(x)^{2}}dx\right)=3\sum_{i=1,b_{i}}^{2g+1}(b_{i},0)-({\widetilde{x}},{\widetilde{y}})-({\widetilde{x}},-{\widetilde{y}})-2\sum_{i=1}^{g}(x_{i},y_{i})-2\sum_{i=1}^{g}(x_{i},-y_{i})-3\infty.$
  3. (3)
    By considering the local parameter $t$ ($x=1/t^{2}$) at $\infty$, we have
    $\mathrm{res}_{\infty}\frac{f(x)}{(x-{\widetilde{x}})F(x)^{2}}dx=-2(\lambda_{2g}+{\widetilde{x}}+2\wp_{gg}).$
  4. (4)
    By considering the local parameter $t$ ($t=x-{\widetilde{x}}$) at $({\widetilde{x}},\pm{\widetilde{y}})$, we have
    $\mathrm{res}_{({\widetilde{x}},\pm{\widetilde{y}})}\frac{f(x)}{(x-{\widetilde{x}})F(x)^{2}}dx=\frac{f({\widetilde{x}})}{F({\widetilde{x}})^{2}}.$
  5. (5)
    By considering the local parameter $t$ ($t=x-x_{k}$) at $(x_{k},\pm y_{k})$, we have
    $\mathrm{res}_{(x_{k},\pm y_{k})}\frac{f(x)}{(x-{\widetilde{x}})F(x)^{2}}dx=\frac{1}{F^{\prime}(x_{k})}\left[\frac{\partial}{\partial x}\left(\frac{f(x)}{(x-{\widetilde{x}})F^{\prime}(x)}\right)\right]_{x=x_{k}}.$
Using them we prove the relation.
Here we introduce ${{\widetilde{\zeta}}_{i}(P):=-\int^{P}_{\infty}{\nu^{\mathrm{II}}_{i}}}$.
Lemma 4.11.
Let $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\ldots,\iota_{X}(x_{g},y_{g}))$, $v={\widetilde{w}}(\iota_{X}(x,y))$.Then we have
(4.6) $[\zeta_{g}(u+v)-\zeta_{g}(u-v)-2{\widetilde{\zeta}}_{g}(v)]=\frac{2y}{F(x)}.$
Proof.
The sketch of the proof is as follows. The right hand is the meromorphic function on $J_{X}\times{\widetilde{w}}(\iota_{X}X)$ because of the translational formula (3.9). Thus we show the equality of both sides. Let us show the right hand side is a meromorphic function of $S^{g}X\times X$. We note that both sides are odd function with respect to $v$. For $v$, we expand both sides,
$\zeta_{i}(u+v)-\zeta_{i}(u-v)=-2\sum_{ij}\wp_{ij}v_{j}-\frac{2}{3!}\sum_{ij}\wp_{ijk\ell}v_{j}v_{k}v_{\ell}+d_{>3}(v_{i}).$
The local parameter $t=x^{g+1}/y$ shows them.
Remark 4.12.
Lemma 4.11 is generalized as follows.
(4.7) $\begin{split}&[\zeta_{g-1}(u+v)-\zeta_{g-1}(u-v)+2{\widetilde{\zeta}}_{g-1}(v)]=\frac{2y}{F(x)}(x-\wp_{gg}),\\ &[\zeta_{g-2}(u+v)-\zeta_{g-2}(u-v)+2{\widetilde{\zeta}}_{g-2}(v)]=\frac{2y}{F(x)}(x^{2}-\wp_{gg}x+\wp_{g,g-1}).\end{split}$
Using above relations, we have the following lemma:
Lemma 4.13.
For $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\ldots,\iota_{X}(x_{g},y_{g}))$ and $v={\widetilde{w}}(\iota_{X}(x,y))$
$\zeta_{g}(u+v)-\zeta_{g}(u)-{\widetilde{\zeta}}_{g}(v)=\frac{y}{F(x)}-\sum_{r=1}^{g}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})},$
$\wp_{gg}(u+v)-\wp_{gg}(u-v)=-\sum_{r=1}^{g}\frac{4yy_{r}}{(x-x_{r})F^{\prime}(x_{r})F(x)},$
$\wp_{gg}(u+v)+\wp_{gg}(u)+x=\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}-\frac{y}{F(x)}\right]^{2}-\lambda_{2g}.$
5. Baker function for $X$
Let us generalize (2.11) and $\psi(u;v)=\displaystyle{\frac{{\mathrm{e}}^{-\zeta(v)u}\sigma(u+v)}{\sigma(u)\sigma(v)}}$ of genus one to general $g$.
Let us define
(5.1) $\psi(u;v)=\frac{{\mathrm{e}}^{-\sum_{i}{\widetilde{\zeta}}_{i}(P)u_{i}}\sigma(u+v)}{\sigma(u)\sigma_{\natural_{1}}(v)}$
Here we let ${{\widetilde{\zeta}}_{i}(P):=-\int^{P}_{\infty}{\nu^{\mathrm{II}}_{i}}}$ for $u\in{\mathbb{C}}^{g}$ and $P\in X$, we let $v=w(P)$.
Proposition 4.8, Lemmas 4.11 and 4.13 show the generalization of (2.11).
Proposition 5.1.
For $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\ldots,\iota_{X}(x_{g},y_{g}))$ and $v={\widetilde{w}}(\iota_{X}({\widetilde{x}},{\widetilde{y}}))$, we have
(5.2) $\left[\frac{\partial^{2}}{\partial u_{g}^{2}}-2\wp_{gg}(u)\right]\psi(u;v)=({\widetilde{x}}+\lambda_{2g})\psi(u;v),$
(5.3) $\left[\frac{\partial}{\partial u_{g}}-(\zeta_{g}(u+v)-\zeta_{g}(u))\right]\psi(u;v)={\widetilde{\zeta}}_{g}(v)\psi(u;v).$
Proof.
The same as the case of $g=1$.
6. The algebraic differential relation of genera two and three
We extend the algebraic differential relation (1.1) of genus one to genera two and three.
6.1. Genus two
Let us consider only the genus $g=2$ case as follows. We assume $(P_{1},P_{2},{\widetilde{P}})\in S^{2}X\times X$, $(P_{i}=(x_{i},y_{i})$, and ${\widetilde{P}}=({\widetilde{x}},{\widetilde{y}})$ such that $u={\widetilde{w}}(\iota_{X}(P_{1},P_{2}))$ and $v={\widetilde{w}}(\iota_{X}({\widetilde{P}}))$.
Corollary 4.6 of $g=2$ shows
(6.1) $\frac{\sigma(u+v)\sigma(u-v)}{\sigma(u)^{2}\sigma_{\natural_{1}}(v)^{2}}={\widetilde{x}}^{2}-\wp_{22}{\widetilde{x}}-\wp_{21}.$
We immediately obtain the following relation from (6.1), though we don’t use it in this paper,
Lemma 6.1.
Let us consider the case ${\widetilde{P}}^{(\pm)}=(\pm{\widetilde{x}},{\widetilde{y}}^{(\pm)})$ and $v^{(\pm)}:=w(P^{(\pm)})$. Then we have
(6.2) $\frac{\sigma(u+v^{(+)})\sigma(u-v^{(+)})}{\sigma(u)^{2}\sigma_{\natural_{1}}(v^{(+)})^{2}}-\frac{\sigma(u+v^{(-)})\sigma(u-v^{(-)})}{\sigma(u)^{2}\sigma_{\natural_{1}}(v^{(-)})^{2}}=-2\wp_{22}{\widetilde{x}}.$
Now we state our main theorem in this paper, in which we obtain the generalization of (2.14) or (1.1).
Theorem 6.2.
For $(P_{1},P_{2},{\widetilde{P}})\in S^{2}X\times X$ such that 1) $x_{1}x_{2}=-{\widetilde{x}}^{2}$, or 2) $x_{1}x_{2}=0$, and $\alpha\in{\mathbb{C}}^{g}$ and $c,\beta\in{\mathbb{C}}$, we let
$\Psi(u;v,t):=\frac{1}{\sqrt{{\widetilde{x}}}}{\mathrm{e}}^{{\,{}^{t}\negthinspace}\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u;v,t):=\frac{-1}{\sqrt{{\widetilde{x}}}}{\mathrm{e}}^{-{\,{}^{t}\negthinspace}\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v).$
The following formulae hold.
$\begin{split}\mbox{1)\ :\ }&\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi,\\ \mbox{2)\ :\ }&\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta-{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi.\\ \end{split}$
Proof.
From (5.2), we have
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-\wp_{gg}\Psi=\frac{1}{2}(2\beta+{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi.$
Under the restriction 1), $(\Psi^{*}\Psi)$ is reduced to $(1/{\widetilde{x}})$ times ${\widetilde{x}}\wp_{22}(u)$. whereas under the restriction 2), $(\Psi^{*}\Psi)$ is equal to ${\widetilde{x}}-\wp_{22}(u)$.
Remark 6.3.
As mentioned in Introduction the Baker equation (5.2) contains the potential $\wp_{gg}$, while the terms Weierstrass-$\wp$ in $F({\widetilde{x}})$ in the addition theorem (4.2) do not coincide with $\wp_{gg}$ except for $g=1$. Thus, the discrepancy prevents us from obtaining the generalization of (1.1) directly for a long time.
Even for genus two, we find it by restriction the configurations of points in $S^{2}X$. In other words, the restriction may be effective for the hyperelliptic solution of the NLS equation in terms of $\sigma$ function and enable us to find the generalization to general genus.
However we should consider how the restriction effects on the Jacobian $J_{X}$. Further we note that in the NLS equation, we have crucial problems related to the reality conditions depending on the ”nonfocusing” and ”focusing” NLS equations [P] or the choice of $(\alpha,u)$ which belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$.. However, since these problems are basically complicated as in [Ma4], we will consider these problems as the next steps. we have considered the similar problem for the MKdV equation,
6.2. Genus three
Let us consider the genus $g=3$ case as follows. We assume $(P_{1},P_{2},P_{3}{\widetilde{P}})\in S^{3}X\times X$, $(P_{i}=(x_{i},y_{i})$, and ${\widetilde{P}}=({\widetilde{x}},{\widetilde{y}})$ such that $u={\widetilde{w}}(\iota_{X}(P_{1},P_{2},P_{3}))$ and $v={\widetilde{w}}(\iota_{X}({\widetilde{P}}))$.
Corollary 4.6 of $g=3$ shows
(6.3) $\frac{\sigma(u+v)\sigma(u-v)}{\sigma(u)^{2}\sigma_{\natural_{1}}(v)^{2}}={\widetilde{x}}^{3}-\wp_{33}{\widetilde{x}}^{2}-\wp_{32}{\widetilde{x}}-\wp_{31}.$
Now we state our main theorem in this paper, in which we obtain the generalization of (2.14) or (1.1).
Theorem 6.4.
For $(P_{1},P_{2},P_{3},{\widetilde{P}})\in S^{3}X\times X$ such that ${\widetilde{x}}(x_{1}x_{2}+x_{2}x_{3}+x_{3}x_{1})-x_{1}x_{2}x_{3}=0$, $\alpha\in{\mathbb{C}}^{g}$ and $c,\beta\in{\mathbb{C}}$, we let
$\Psi(u;v,t):=\frac{1}{{\widetilde{x}}}{\mathrm{e}}^{{\,{}^{t}\negthinspace}\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u;v,t):=\frac{-1}{{\widetilde{x}}}{\mathrm{e}}^{-{\,{}^{t}\negthinspace}\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v).$
The following formula holds.
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta-{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi,$
where $\wp_{33}={\widetilde{x}}(\gamma_{1}+\gamma_{2}+1/\gamma_{1}\gamma_{2})$.
Proof.
From (5.2), we have
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-\wp_{gg}\Psi=\frac{1}{2}(2\beta+{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi.$
Under the restriction, $(\Psi^{*}\Psi)$ is reduced to $(1/{\widetilde{x}})^{2}$ times ${\widetilde{x}}^{3}-{\widetilde{x}}^{2}\wp_{33}(u)$.
Acknowledgments:
The author is grateful to Professor Emma Previato for suggestions on this study at a very early stage and to Dr. Kayo Kinjo for discussing the algebraic solution of the NLS equation based on [BBEIM]. The author has been supported by the Grant-in-Aid for Scientific Research (C) of Japan Society for the Promotion of Science Grant, No.21K03289.
Competing interests:
The author declares none.